OpenStudy (lina777):

Help please! Mark all statements that are true: A. If the measure of theta is 60 degrees, then the arc length is r. B. If the measure of theta is .5 radians, then the arc length is r/2. C. The ratio of the arc length to r is always equal to 1/pi. D. If the ratio fo the arc length to r is 1, then the measure of theta is 1 radian.

jimthompson5910 (jim_thompson5910):

for part A, use this formula \[\Large s = \frac{x}{360}*2*\pi*r\] s = arc length x = angle in degrees r = radius

jimthompson5910 (jim_thompson5910):

In this case, x = 60

jimthompson5910 (jim_thompson5910):

for part A they're asking if x = 60, does plugging this in yield s = r ?

OpenStudy (lina777):

Okay! Thank you! But how does that help me to know if s=r?

jimthompson5910 (jim_thompson5910):

what do you get when you plug in x = 60 and simplify

OpenStudy (lina777):

1/3 pi x r=s

jimthompson5910 (jim_thompson5910):

so the result we get is NOT just 'r' all by itself so the claim made in choice A is false

OpenStudy (lina777):

Oh I see! That's great, thank you very much!

jimthompson5910 (jim_thompson5910):

IF it said that the angle was `1 radian` then the claim would be true if the angle was `1 radian` then the arc length would be equal to the radius r formula s = theta*r where s = arc length theta is in radians (this is very important) r = radius

OpenStudy (lina777):

Hm. So, which part of the question does that help with? Sorry, this is just a trickily worded question.

jimthompson5910 (jim_thompson5910):

that's just extra info, but it will help out with B

OpenStudy (lina777):

Great. Thank you!

jimthompson5910 (jim_thompson5910):

hint: 0.5 = 1/2

jimthompson5910 (jim_thompson5910):

s = theta*r for part B, theta = 1/2 radians

OpenStudy (lina777):

So B would be true then?

jimthompson5910 (jim_thompson5910):

yes because s = theta*r s = (1/2)*r s = r/2

OpenStudy (lina777):

Great. So, I didn't think C was correct. Am I right in my thinking or not?

jimthompson5910 (jim_thompson5910):

why are you thinking that

OpenStudy (lina777):

I don't know, just that it seems as if there is not a distinct correlation. Ha. But I don't really know :)

jimthompson5910 (jim_thompson5910):

what do you get when you compute `s/r` replace the 's' with the right side of the first formula I posted

jimthompson5910 (jim_thompson5910):

what terms do you see cancelling out?

OpenStudy (lina777):

the radius?

jimthompson5910 (jim_thompson5910):

yes r cancels leaving you with `(2pi*x)/360 = (pi*x)/180` this is NOT equal to `1/pi`

OpenStudy (lina777):

Great! My intuition was correct, thanks for solidifying that! Okay, for the last one, I put that that was correct. Is that true?

OpenStudy (lina777):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's find out

jimthompson5910 (jim_thompson5910):

dividing s/r using the second formula given yields this s/r = (theta*r)/r = theta so the ratio of the arc length s to the radius r is equal to theta this only true if theta is in radians

jimthompson5910 (jim_thompson5910):

if the ratio is equal to 1, then theta is equal to 1 radian

jimthompson5910 (jim_thompson5910):

so yep D is true

OpenStudy (lina777):

You are a lifesaver! Thanks so very much!

OpenStudy (lina777):

I got the question right, thank you! @jim_thompson5910